Write the balanced overall equation for the followingprocess (equation 3) and calculate

∆HoveralI.


(1) N2 (g) + O2(g) → 2NO(g) ∆H = 180.6 kJ

(2) 2NO(g) + O2(g) → 2NO2(g) ∆H = - 114.2 kJ

(3) ∆Hoverall=?


1
Expert's answer
2021-09-27T06:33:35-0400

Reaction

"N_2 (g) + O_2(g) \u2192 2NO(g)"

"NO_2 (g) \to NO(g) +\frac{1}{2}O_2(g)"

"N_2O (g) \to N_2(g)+ \frac{1}{2}O_2(g)"


"N_2O (g) + NO_2(g) \u2192 3NO(g)"


"∆H=∆H_1+(-\frac{1}{2}∆H_2)+\frac{1}{2}∆H_3"


"∆H=180.6+(-\frac{1}{2}-114.2)+\frac{1}{2}(-163.2)"


"\u2206H=156.1kJ"